# CNSocialDepress: A Chinese Social Media Dataset for Depression Risk Detection and Structured Analysis

### v1

@@ Appendix A Appendix
A.1 Baseline Models
  • DeepSeek-R1 is a large-scale model specialized in logical reasoning,This model excels in complex reasoning tasks such as mathematics and programming while maintaining strong overall performance on general tasks. Chosen for its specialized excellence in complex logical reasoning while maintaining strong general task performance, potentially beneficial for in-depth structured psychological analysis and nuanced depression risk detection.
  • DeepSeek R1 Distill-14B is a lightweight reasoning model derived from Qwen2.5-14B through Knowledge Distillation. Chosen for its lightweight reasoning capabilities, aiming to retain the advantages of larger models in complex tasks while reducing computational cost, making it suitable for efficient depression risk analysis.
  • Qwen2.5-14B is part of Alibaba’s Qwen series of large language models. Chosen for its strong performance in long-form text generation and instruction-following, relevant for generating structured psychological analyses from Chinese social media posts.
  • GPT-4o is a multimodal large language model released by OpenAI. Chosen for its state-of-the-art multimodal and multilingual processing capabilities, including Chinese, and its improved overall performance and response speed, making it a powerful baseline for depression detection and psychological analysis.
  • GPT-4o-Mini is a lightweight AI model developed by OpenAI to provide an efficient and cost-effective AI solution. Chosen as a resource- and cost-efficient model with robust text understanding and generation, suitable for less computationally demanding depression risk identification and basic psychological analysis tasks.
  • Llama3-8B-Chinese-Chat, derived from Meta-Llama-3-8B, is an instruction-tuned model optimized for Chinese dialogue tasks. It has been fine-tuned to process both Chinese and English inputs. Chosen for its optimization for Chinese dialogue tasks and ability to handle extended conversations, making it well-suited for analyzing conversational social media posts for depression risk and generating coherent psychological insights in Chinese.
A.2 20 Randomly Generated Expression for the Positive Users
@@ A.1 Baseline Models
A.3 20 Distinct Label Expressions for the Negative Users
@@
A.4 Instructions for the Fine-Tuning Process
@@
A.5 Module II Prompt
@@
A.6 Prompt for Table 3 On Section 5.3.1
Prompt is as follows: You are a professional in psychology and text sentiment analysis. Please determine whether the user is at risk of depression based on the following six dimensions:
  • Negative emotions (secondary criterion)
  • Depressive psychological state (primary criterion)
  • Depression-related clinical symptoms (primary criterion)
  • Potential external factors causing depression (secondary criterion)
  • Depression-related medical expressions (primary criterion)
  • Depression-related language expression patterns (secondary criterion)
For each dimension, please cite the corresponding text reference and provide a concise explanation.
[Output Data Example]
  • Answer: “Yes” or “No.”
  • Briefly explain the manifestations in each of the six dimensions, and cite the corresponding text reference.
  • In conclusion, “The user is at risk of depression” or “The user is not at risk of depression.”
A.7 Bert type models
BERT-base-chinese [^9] https://huggingface.co/google-BERT/BERT-base-chinese is built on the BERT-Base architecture Devlin et al. (2019) and is pre-trained on extensive Chinese corpora. It adopts masked language modeling (MLM) and next sentence prediction (NSP) to learn deep bidirectional representations of Chinese text at the character level.
Chinese-RoBERTa-wwm-ext Cui et al. (2019) leverages RoBERTa Liu et al. (2019) with a whole word masking (WWM) approach. Instead of masking individual characters, it masks entire words during MLM, leading to a richer understanding of context. The extended pre-training (ext) on large Chinese corpora further refines its effectiveness in various NLP tasks.
StructBERT-mental Wang et al. (2019) is based on StructBERT, which incorporates linguistic structures into its pre-training. By focusing on word and sentence structure, and adding specific pre-training for mental health texts, it improves performance in identifying and categorizing mental health-related content.

### v2

@@ Appendix A Appendix
A.1.  Baseline Models
  • DeepSeek-R1 is a large-scale model designed for complex reasoning. It performs well on tasks such as mathematics and programming, while remaining competitive on general-purpose benchmarks. We include it as a strong reasoning-oriented baseline for structured psychological analysis and depression risk assessment.
  • DeepSeek-R1-Distill-14B is a distilled reasoning model derived from Qwen2.5-14B via knowledge distillation. It aims to retain key reasoning capabilities at lower computational cost. We include it as an efficient baseline for depression risk analysis.
  • Qwen2.5-14B is a large language model from the Qwen series. It provides strong instruction following and long-form generation. We include it as a Chinese-capable baseline for generating structured psychological analyses from social media posts.
  • GPT-4o is a multimodal large language model released by OpenAI. It supports multilingual generation, including Chinese. We include it as a strong general-purpose baseline for depression detection and psychological analysis.
  • GPT-4o-mini is a smaller model in the GPT-4o family. It is designed for efficiency and lower inference cost. We include it as a lightweight baseline for depression risk identification and structured summarization.
  • Llama3-8B-Chinese-Chat, derived from Meta-Llama-3-8B, is an instruction-tuned model for Chinese dialogue. It supports both Chinese and English inputs. We include it as a Chinese dialogue-oriented baseline for user-level depression risk analysis and generation.
A.2.  20 Randomly Generated Expression for the Positive Users
@@ A.1.  Baseline Models
A.3.  20 Distinct Label Expressions for the Negative Users
@@
A.4.  Instructions for the Fine-Tuning Process
@@
A.5.  BERT-based Models
BERT-base-chinese [^8] https://huggingface.co/google-bert/bert-base-chinese follows the BERT-Base architecture Devlin et al. (2019) and is pre-trained on large-scale Chinese corpora. It uses masked language modeling (MLM) and next sentence prediction (NSP). The model learns bidirectional representations for Chinese text, typically at the character level.
Chinese-RoBERTa-wwm-ext Cui et al. (2019) is based on RoBERTa Liu et al. (2019b) and uses whole word masking (WWM). It masks complete words rather than individual characters during MLM. This design can better capture word-level semantics in Chinese. The extended pre-training (“ext”) on large Chinese corpora further improves robustness across NLP tasks.
StructBERT-mental Wang et al. (2019) builds on StructBERT, which incorporates structural objectives during pre-training. It emphasizes word- and sentence-level structure. In addition, it is adapted for mental health text, which can benefit mental health-related classification.
A.6.  Classification Performance Comparison
Model/Test Dataset Swdd Origin Wu3d CNSD (Gold)
Acc F1 Acc F1 Acc F1
RoBERTa Chinese (trained by swdd origin) 0.89 0.90 0.76 0.80 0.88 0.87
RoBERTa Chinese (trained by wu3d) 0.88 0.87 0.90 0.91 0.82 0.78
Bert based Chinese (trained by wu3d) 0.88 0.87 0.90 0.91 0.83 0.80
StructBERT-mental (trained by wu3d) 0.85 0.85 0.92 0.92 0.83 0.80
Llama3-8b 0.86 0.86 0.79 0.79 0.60 0.63
Glm4-9b-chat 0.76 0.81 0.77 0.81 0.79 0.82
Qwen2.5 7b 0.88 0.89 0.84 0.85 0.94 0.94
GPT-4o-Mini 0.65 0.74 0.65 0.74 0.53 0.68
GPT-4o 0.92 0.92 0.91 0.92 0.83 0.86
DeepSeek-R1 671b 0.86 0.87 0.89 0.90 0.81 0.84
A.7.  Module II Prompt